OpenStudy (anonymous):

(k^2-14)/(8k^2+3k) divided by (k-2)/(16k+6) simplify. My work: (k^2-14)/(8k^2+3k) divided by (k-2)/(16k+6) (k^2-14)/(8k^2+3k) times (16k+6)/(k-2) I crossed out like terms and got: (k+2)/(k(k+1)) * 2k+2 Am I correct or no??

OpenStudy (anonymous):

@satellite73 Please help?

OpenStudy (anonymous):

@timo86m please help??

OpenStudy (anonymous):

((k^2-14)/(8k^2+3k)) /( (k-2)/(16k+6)) that?

OpenStudy (jim766):

\[\frac{ k^2-14 }{ 8k^2+3k }\times \frac{ 16k + 6 }{ k-2 } \] now lets factor what we can \[\frac{ k^2-14 }{ k(8k+3) }\times \frac{ 2(8k + 3) }{ k-2 } \] now we can cancel out \[\frac{ k^2-14 }{ k }\times \frac{ 2 }{ k-2 } \]

OpenStudy (anonymous):

and it seems you did the first thing right of multiply by reciprocal.

OpenStudy (jim766):

unless there is a typo and the k^2 -14 is really k^2 -4 all I can see to do now is multiply across

OpenStudy (anonymous):

(2 k^2-28)/(k^2-2 k)

OpenStudy (anonymous):

i'd go with jim

OpenStudy (jim766):

makes sense?

OpenStudy (anonymous):

Not really. On the first part, how do you get just k in the denominator??

OpenStudy (anonymous):

Oh wait, nevermind

OpenStudy (anonymous):

Oh and it was a typo.It is 4 not 14. Dx

OpenStudy (anonymous):

(k^2-4)/(8k^2+3k) divided by (k-2)/(16k+6) Thats what it SHOULD be. @Jim766

OpenStudy (jim766):

\[\frac{ k^2-4 }{ k}\times \frac{ 2 }{ k-2 }\] now the k^2 -4 factors because it is a difference of 2 squares \[\frac{ (k +2)(k - 2) }{ k}\times \frac{ 2 }{ k-2 }\] now the k-2 cancels out

OpenStudy (jim766):

\[\frac{ (k +2) }{ k}\times \frac{ 2 }{ 1 }\] now mult straight across \[\frac{ 2k +4 }{ k}\] and that is what you have left

OpenStudy (jim766):

when you do these types of problems, if it is division, change it to multiplication like you did. once that is done look for ways to factor the different parts. Very often when everything is factored, you can see what has to be cancelled out. What is left is you answer. questions?

OpenStudy (anonymous):

Nope, makes sense. Thank you

OpenStudy (jim766):

yw